# Impurity dynamics in a zero-temperature gas

### v1

@@ IV Position and momentum distributions
[…]
Let us ignore correlations and try to probe the position distribution using a physically appealing but uncontrolled approximation. Employing a diffusion equation with a time-dependent diffusion coefficient is a naive approach. Indeed, the hydrodynamic radial velocity leads to $H\sim R_{\text{imp}}$ drift comparable with displacement generated by diffusion. A more consistent but still uncontrolled approximation is based on a convection-diffusion equation. Focusing on time dependence, ignoring numerical factors, and taking into account the spherical symmetry, we get a convection-diffusion equation
$\big{(}\partial_{t}+t^{-\frac{113}{175}}\partial_{r}\big{)}P=t^{-\frac{51}{175}}\big{(}\partial_{r}^{2}+\tfrac{2}{r}\partial_{r}\big{)}P$ (24)
in three dimensions. We have also taken into account the spherical symmetry. We seek the solution of the convection-diffusion equation (24) in the scaling form
$P(r,t)=t^{-\frac{186}{175}}\,\mathcal{P}(\eta),\qquad\eta=t^{-\frac{62}{175}}r$ (25)
@@
[…]
Plugging (25) into (24) we obtain a second-order ordinary differential equation
$-\frac{186}{175}\,\mathcal{P}-\frac{62}{175}\,\eta\,\mathcal{P}^{\prime}=\mathcal{P}^{\prime\prime}+\frac{2}{\eta}\,\mathcal{P}^{\prime}$ (27)
where prime denotes the derivative with respect to $\eta$. Multiplying Eq. (27) by $\eta^{2}$ and integrating yields the first-order ordinary differential equation
$-\frac{62}{175}\,\eta^{3}\mathcal{P}=\eta^{2}\mathcal{P}^{\prime}$ (28)
The solution to Eq. (28) satisfying the normalization requirement (26) is
$\mathcal{P}=\left(\frac{31}{175\pi}\right)^{\frac{3}{2}}\exp\!\left[-\frac{31}{175}\,\eta^{2}\right]$ (29)
The omitted numerical factors do not affect the Gaussian form. Thus, the convection-diffusion approach leads to the Gaussian solution in three dimensions. The extension to any $d\geq 2$ is straightforward and gives the same result:
$P(\bm{r},t)=\frac{1}{(2\pi R_{\text{imp}}^{2})^{d/2}}\,\exp\!\left[-\frac{r^{2}}{2R_{\text{imp}}^{2}}\right]$ (30)
with $R_{\text{imp}}^{2}=C_{d}\lambda^{2}(R/\lambda)^{2h_{d}}$. The numerical factor $C_{d}$ cannot be fixed in the realm of the above approach. We emphasize that have not justified the convection-diffusion approach, so the Gaussian form (30) of the position distribution is questionable.
@@
$(\partial_{t}+\bm{v}\cdot\nabla)\Pi(\bm{r},\bm{v},t)=4\pi a^{2}[\Gamma_{+}-\Gamma_{-}]$ (31a)
@@
$\bm{v}^{*}=\bm{v}+\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})],\quad\bm{w}^{*}=\bm{w}-\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})]$ (32)
[…]
$F(\bm{r},\bm{w},t)=\frac{\rho(r,t)}{[2\pi T(r,t)]^{3/2}}\,\exp\!\left\{-\frac{m|\bm{w}-u(r,t)|^{2}}{2T(r,t)}\right\}$
The density $\rho(r,t)$, temperature $T(r,t)$, and radial velocity $u(r,t)$ satisfy the Navier-Stokes equations in the core region. In the long time limit, these scalar fields acquire the scaling form depending on a single variable $r/H$. The scaled density, temperature, and radial velocity satisfy non-linear coupled ordinary differential equations that have not been solved in the core region.
@@
Another challenge is to analyze the number of collisions experienced by the impurity beyond the typical value, Eq. (9), estimated via arguments based on kinetic theory. Even for the dilute hard-sphere gas in thermal equilibrium, the collisional statistics of the tagged particle is remarkably subtle Lue (2005); Visco et al. (2008a, b); Paik (2014).
@@ V Numerical verification
To test the predictions about the characteristics of impurities, we simulated hard discs in two dimensions using an event-driven molecular dynamics (EDMD) scheme. The system is prepared by randomly distributing discs, while ensuring the no overlap condition, within a simulation box of dimensions $L\times L$, at a fixed specified number density $n_{\infty}=N/L^{2}$ and volume fraction ($\phi=\pi a^{2}n_{\infty}$). We chose four discs closest to the origin and gave them random momenta, ensuring that the total energy is exactly $E$ and the total momentum vanishes. The simulation proceeds by calculating the times of all possible future collisions between discs and with the walls, then evolving the system to the earliest collision event. At each event, the velocities of the colliding particles are updated based on the conservation of momentum and energy. Between collisions, the discs move linearly with constant velocity. In all simulations performed, we used density $n_{\infty}=0.15$, the radius of particles $a=1/2$, and the mass of particles $m=1$.
The system contained $N=40000$ particles, and the EDMD simulations ran until $t=4000$. For computing the properties of impurities, we averaged over $30000$ randomly chosen initial conditions, while for computing bulk and surface quantities, we used $5000$ initial conditions. For any observable, say $A$, we expect power law behavior, $A\sim t^{\alpha}$, in the long time limit, while at finite times, there will be corrections. Hence, in our numerics, we estimate the temporal evolution of this exponent for various quantities of interest. We recorded observables at exponentially separated times in powers of $2^{1/4}$. We then obtain the exponents from the data at successive times (namely $\alpha(t)=4[\ln A(2^{1/4}t)-\ln A(t)]/\ln 2$).
$\bm{R}_{imp}$ $\bm{C}_{imp}$ $\bm{V}_{imp}$ $\bm{u}_{imp}$ $\bm{T}_{0}$ $\bm{n}_{0}$
$\alpha_{T}$ $2/5$ $2/5$ $-2/5$ $-3/5$ $-4/5$ $-1/5$
$\alpha_{S}$ $0.423$ $0.363$ $-0.399$ $-0.446$ $-0.692$ $-0.168$
@@
Table 2: The table shows the comparison between theoretically computed and numerically estimated exponents $\alpha$ for time $(t)$ dependencies of the form $t^{\alpha}$ for the quantities of interests. $\alpha_{T}$ and $\alpha_{S}$ stands for Theory and Simulation.
@@
where the average is over $n_{R}=3\times 10^{4}$ different initial conditions and over the four impurities. We define $C_{imp}(t)$ as the mean number of collisions that any impurity undergoes up to time $t$. The comparison between the kinetic predictions and the simulation results are shown in Fig. 2$(A)$-$(C)$.
To compute the fluid speed, $u_{imp}$, around the impurity particle, we find the average flow velocity of all the particles within a fixed radius around the impurity and compute its magnitude. The fixed radius is chosen so that, at $t=0$, there are $100$ other gas particles within that distance. Again, we take an average over the four impurities and $n_{R}$ realizations. The comparison of $u_{\text{imp}}$ between theory and simulations is shown in Fig. 2$(D)$.
In Fig. 3 we plot the full probability distribution function of the impurity’s displacement at time $t$. From Eq. 30, the predicted form is
@@
Our simulations find close agreement with this.
Figure 3: Scaled probability distribution function for impurity displacement plotted at different times. The dashed curve in the figure is plotted by fitting Eq. (34) with $R_{\text{imp}}=b~{}t^{2/5}$ against the PDF obtained from simulation data at $t=4000$. The parameter $b$ obtained by fitting is $1.75$ which is close to $b=1.71$, obtained directly from computation of $R_{imp}/t^{2/5}$ at $t=4000$.
@@ VI Summary
Impurities are generally different from the host particles. The disparity between the mass and radius of impurity and the mass and radii of host spheres may lead to interesting behaviors, especially when these disparities are significant. We leave this problem for the future. One can also explore the impurity problem in high spatial dimensions, $d\gg 1$, which may be more amenable to analytical treatment and perhaps sheds light on the behavior in the physically relevant three-dimensional case. Finally, we mention a different setting where the impurity is normally incident on the static gas occupying the half-space Antal et al. (2008). The impurity, as well as every host sphere, is expected to be ejected backward into an initially empty half-space after a finite number of collisions. This assertion is intuitively clear, albeit proof could be challenging. So far, the splash problem has been quantitatively studied in one-dimension Chakraborti et al. (2022); Kumar and Rajesh (2025b).

### v2

@@ IV Position and momentum distributions
$P(\bm{r},t)=\int d\bm{v}\,\Pi(\bm{r},\bm{v},t),\quad Q(\bm{v},t)=\int d\bm{r}\,\Pi(\bm{r},\bm{v},t)$
Let us ignore correlations and try to probe the position distribution using a physically appealing but uncontrolled approximation. Employing a diffusion equation with a time-dependent diffusion coefficient is a naive approach. Indeed, the hydrodynamic radial velocity leads to $H\sim R_{\text{imp}}$ drift comparable with displacement generated by diffusion. A more consistent but still uncontrolled approximation is based on a convection-diffusion equation,
$(\partial_{t}+u_{\rm imp}\partial_{r})P=D_{0}\nabla^{2}P.$ (25)
For two-dimensional hard discs gas, we have $u_{\rm imp}\sim t^{-3/5}$ while $D_{0}\sim t^{-1/5}$. Focusing on time dependence, ignoring numerical factors, and taking into account the spherical symmetry, we get a convection-diffusion equation, which in two-dimensions is given by:
$\big(\partial_{t}+t^{-\frac{3}{5}}\partial_{r}\big)P=t^{-\frac{1}{5}}\big(\partial_{r}^{2}+\tfrac{1}{r}\partial_{r}\big)P.$ (26)
We seek the solution of Eq. (26) in the scaling form
$P(r,t)=t^{-\frac{4}{5}}\,\mathcal{P}(\eta),\qquad\eta=t^{-\frac{2}{5}}r.$ (27)
@@
$\int_{0}^{\infty}dr\,2\pi r\,P(r,t)=2\pi\int_{0}^{\infty}d\eta\,\eta\mathcal{P}(\eta)=1.$ (28)
Plugging Eq. (27) into Eq. (26) we obtain a second-order ordinary differential equation,
$\frac{4}{5}\,\mathcal{P}-\frac{2}{5}\,\eta\,\mathcal{P}^{\prime}=\mathcal{P}^{\prime\prime}+\frac{1}{\eta}\,\mathcal{P}^{\prime},$ (29)
where prime denotes the derivative with respect to $\eta$. Multiplying Eq. (29) by $\eta$ and integrating yields the first-order ordinary differential equation,
$\frac{2}{5}\,\eta^{2}\mathcal{P}=\eta\mathcal{P}^{\prime}.$ (30)
The solution to Eq. (30) satisfying the normalization requirement Eq. (28) is
$\mathcal{P}=\left(\frac{1}{5\pi}\right)\exp\!\left[-\frac{1}{5}\,\eta^{2}\right].$ (31)
The same approach can be repeated in three dimensions as well to obtain similar result in $3$-dimensions. In 3D, the analog of Eq. (26) is,
$\big(\partial_{t}+t^{-\frac{113}{175}}\partial_{r}\big)P=t^{-\frac{51}{175}}\big(\partial_{r}^{2}+\tfrac{2}{r}\partial_{r}\big)P.$ (32)
while the analog of the scaling form Eq. (27) is,
$P(r,t)=t^{-\frac{186}{175}}\,\mathcal{P}(\eta),\qquad\eta=t^{-\frac{62}{175}}r.$ (33)
Plugging Eq. (33) into Eq. (32) and solving the resulting ODE gives,
$\mathcal{P}=\left(\frac{31}{175\pi}\right)^{\frac{3}{2}}\exp\!\left[-\frac{31}{175}\,\eta^{2}\right].$ (34)
The omitted numerical factors do not affect the Gaussian form. Thus, the convection-diffusion approach leads to the Gaussian solution in two-dimensions. The extension to any $d\geq 2$ is straightforward and gives the same result:
$P(\bm{r},t)=\frac{1}{(2\pi R_{\text{imp}}^{2})^{d/2}}\,\exp\!\left[-\frac{r^{2}}{2R_{\text{imp}}^{2}}\right],$ (35)
with $R_{\text{imp}}=C_{d}\lambda(R/\lambda)^{h_{d}}$. The numerical factor $C_{d}$ cannot be fixed in the realm of the above approach. We emphasize that have not justified the convection-diffusion approach, so the Gaussian form (35) of the position distribution is questionable.
@@
$(\partial_{t}+\bm{v}\cdot\nabla)\Pi(\bm{r},\bm{v},t)=4\pi a^{2}[\Gamma_{+}-\Gamma_{-}].$ (36a)
@@
$\bm{v}^{*}=\bm{v}+\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})],\quad\bm{w}^{*}=\bm{w}-\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})],$ (37)
where $\bm{e}\in\mathbb{S}^{2}$ is the deflection angle and the integration measure is normalized: $\int d\bm{e}=1$. Due to local thermodynamic equilibrium, the joint distribution of the host spheres is locally Maxwell
$F(\bm{r},\bm{w},t)=\frac{\rho(r,t)}{[2\pi T(r,t)]^{3/2}}\,\exp\!\left\{-\frac{m|\bm{w}-u(r,t)|^{2}}{2T(r,t)}\right\}.$
The density $\rho(r,t)$, temperature $T(r,t)$, and radial velocity $u(r,t)$ satisfy the Navier-Stokes equations in the core region. In the long time limit, these scalar fields acquire the scaling form depending on a single variable $r/H$. The scaled density, temperature, and radial velocity satisfy nonlinear coupled ordinary differential equations that have not been solved in the core region.
@@
Another challenge is to analyze the number of collisions experienced by the impurity beyond the typical value, Eq. (10), estimated via arguments based on kinetic theory. Even for the dilute hard-sphere gas in thermal equilibrium, the collisional statistics of the tagged particle is remarkably subtle Lue (2005); Visco et al. (2008b, a); Paik (2014).
@@ V Numerical verification
To test the predictions about the characteristics of impurities, we simulated hard discs in two-dimensions using an event-driven molecular dynamics. The system is prepared by randomly distributing discs, while ensuring the no overlap condition, within a simulation box of dimensions $L\times L$, at a fixed specified number density $n_{\infty}=N/L^{2}$ and volume fraction ($\phi=\pi a^{2}n_{\infty}$). We chose four discs closest to the origin and gave them random momenta, ensuring that the total energy is exactly $E$ and the total momentum vanishes. The simulation proceeds by calculating the times of all possible future collisions between discs and with the walls, then evolving the system to the earliest collision event. At each event, the velocities of the colliding particles are updated based on the conservation of momentum and energy. Between collisions, the discs move linearly with constant velocity. In all simulations performed, we used density $n_{\infty}=0.15$, the radius of particles $a=1/2$, and the mass of particles $m=1$.
The system contained $N=40000$ particles, and the event-driven molecular dynamics simulations ran until $t=4000$. For computing the properties of impurities, we averaged over $30000$ randomly chosen initial conditions, while for computing bulk and surface quantities, we used $5000$ initial conditions. For any observable, say $A$, we expect power law behavior, $A\sim t^{\alpha}$, in the long time limit, while at finite times, there will be corrections. Hence, in our numerics, we estimate the temporal evolution of this exponent for various quantities of interest. We recorded observables (averaged over realizations) at exponentially separated times in powers of $2^{1/4}$. We then obtained the running time exponents from the data at successive times (namely $\alpha(t)=4[\ln A(2^{1/4}t)-\ln A(t)]/\ln 2$). In Table (2), we also show the errors in computing these exponents, $\delta\alpha$.
$R_{imp}$ $C_{imp}$ $V_{imp}$ $u_{imp}$ $T_{0}$ $n_{0}$
$\alpha_{T}$ $2/5$ $2/5$ 2/5-2/5 3/5-3/5 4/5-4/5 1/5-1/5
$\alpha_{S}$ $0.423$ $0.363$ 0.399-0.399 0.446-0.446 0.692-0.692 0.168-0.168
@@
Table 2: The table shows the comparison between theoretically computed and numerically estimated exponents $\alpha$ for time $(t)$ dependencies of the form $t^{\alpha}$ for the quantities of interests. $\alpha_{T}$ and $\alpha_{S}$ stands for Theory and Simulation. $\delta\alpha$ is the error in estimation of exponent $\alpha$ due to statistical error in the respective quantity.
@@
where the average is over $n_{R}=3\times 10^{4}$ different initial conditions and over the four impurities. We define $C_{imp}(t)$ as the mean number of collisions that any impurity undergoes up to time $t$. The comparison between the kinetic predictions and the simulation results are shown in Fig. 2(A)(A)-$(C)$.
To compute the fluid speed, $u_{imp}$, around the impurity particle, we find the average flow velocity of all the particles within a fixed radius around the impurity and compute its magnitude. The fixed radius is chosen so that, at $t=0$, there are $100$ other gas particles within that distance. Again, we take an average over the four impurities and $n_{R}$ realizations. The comparison of $u_{\text{imp}}$ between theory and simulations is shown in Fig. 2(D)(D).
Finally we numerically computed the radial distribution function, $p_{\rm imp}(r,t)$, of the impurity’s displacement at time $t$ and compared it with the predicted form from Eq. 35,
@@
In Fig. 3 we see that there is close agreement between the simulation results and the predicted form.
Figure 3: Scaled radial probability distribution function $(p_{\rm imp})$ for impurity (obtained from simulation) plotted at different times. The dashed curve in the figure is plotted by fitting Eq. (39)) with $R_{\text{imp}}=b~t^{2/5}$ against the PDF obtained from simulation data at $t=4000$. The parameter $b$ obtained by fitting is $1.75$ which is close to $b=1.71$, obtained directly from computation of $R_{imp}/t^{2/5}$ at $t=4000$.
@@ VI Summary
Impurities are generally different from the host particles. The disparity between the mass and radius of impurity and the mass and radii of host spheres may lead to interesting behaviors, especially when these disparities are significant. We leave this problem for the future. One can also explore the impurity problem in high spatial dimensions, $d\gg 1$, which may be more amenable to analytical treatment and perhaps sheds light on the behavior in the physically relevant three-dimensional case. Finally, we mention a different setting where the impurity is normally incident on the static gas occupying the half-space Antal et al. (2008). The impurity, as well as every host sphere, is expected to be ejected backward into an initially empty half-space after a finite number of collisions. This assertion is intuitively clear, albeit proof could be challenging. So far, the splash problem has been quantitatively studied in one-dimension Chakraborti et al. (2022); Kumar and Rajesh (2025c); Holovatch et al. (2025).